### v1

@@ 3 The rate of convergence
$\begin{array}[]{rl}0&\geq k(k+1)\|z^{k}-P_{c}(z^{k})\|^{2}+2(k+1)\langle z^{k}-P_{c}(z^{k}),z^{k}-z^{0}\rangle+(C_{1}+C_{2})+C_{3}+\displaystyle\sum_{j=1}^{k}\displaystyle\frac{2j^{2}}{j+1}\|\eta^{j-1}\|^{2}\\[6.0pt] &=k(k+1)\|z^{k}-P_{c}(z^{k})\|^{2}+2(k+1)\langle z^{k}-P_{c}(z^{k}),z^{k}-z^{0}\rangle+D_{1}+D_{2}+E_{1}+E_{2}+\displaystyle\sum_{j=1}^{k}\displaystyle\frac{2j^{2}}{j+1}\|\eta^{j-1}\|^{2}\\[6.0pt] &=k(k+1)\|z^{k}-P_{c}(z^{k})\|^{2}+2(k+1)\langle z^{k}-P_{c}(z^{k}),z^{k}-z^{0}\rangle+(D_{1}+E_{2})+(D_{2}+E_{1})+\displaystyle\sum_{j=1}^{k}\displaystyle\frac{2j^{2}}{j+1}\|\eta^{j-1}\|^{2}\\[6.0pt] &=k(k+1)\|z^{k}-P_{c}(z^{k})\|^{2}+2(k+1)\langle z^{k}-P_{c}(z^{k}),z^{k}-z^{0}\rangle-2k\left\langle z^{k}-z^{0},\eta^{k-1}\right\rangle\\[6.0pt] &\quad-\displaystyle\sum_{j=0}^{k-1}2\left\langle z^{0}-z^{j},\frac{j+1}{j+2}\eta^{j}\right\rangle-\displaystyle\sum_{j=0}^{k-1}2j\left\langle z^{0}-z^{j},\frac{j+1}{j+2}\eta^{j}-\eta^{j-1}\right\rangle\\[6.0pt] &\quad-2k(k+1)\left\langle z^{k}-P_{c}(z^{k}),\eta^{k-1}\right\rangle+2\displaystyle\sum_{j=0}^{k-1}\displaystyle\frac{(j+1)^{3}}{j+2}\left\langle P_{c}(z^{j})-z^{j},\eta^{j-1}-\eta^{j}\right\rangle\\[6.0pt] &\quad-2\displaystyle\sum_{j=0}^{k-1}\displaystyle\frac{j+1}{j+2}\left\langle P_{c}(z^{j})-z^{j},\eta^{j-1}\right\rangle+\displaystyle\sum_{j=1}^{k}\displaystyle\frac{2j^{2}}{j+1}\|\eta^{j-1}\|^{2}.\end{array}$
@@
    $\sum_{j=0}^{k-1}\left\langle z^{0}-z^{j},\frac{j+1}{j+2}\eta^{j}\right\rangle\leq\displaystyle\displaystyle\sum_{j=0}^{k-1}\|z^{0}-z^{j}\|\cdot\|\frac{j+1}{j+2}\eta^{j}\|\leq\kappa_{0}\beta_{0};$
@@
    $\sum_{j=0}^{k-1}\displaystyle\frac{j+1}{j+2}\left\langle P_{c}(z^{j})-z^{j},\eta^{j-1}\right\rangle\leq\displaystyle\sum_{j=0}^{k-1}\displaystyle\frac{j+1}{j+2}\|P_{c}(z^{j})-z^{j}\|\|\eta^{j-1}\|\leq\displaystyle\kappa_{0}\displaystyle\sum_{j=0}^{k-1}\displaystyle\varepsilon_{j-1}\leq\displaystyle\kappa_{0}\beta_{0}.$
@@
Next, we show that with the specific choice of the tolerance, the Halpern’s acceleration proximal point method can achieve the convergence quite close to the tight convergence rate in [6].
@@
Corollary 3.1 provides explicit convergence rate estimates for the Halpern accelerated proximal point method under the tolerance rule (3.20). It shows that the residual $\|z^{k}-P_{c}(z^{k})\|$, which measures the violation of the monotone inclusion, converges to zero at a sublinear rate depending on the decay parameter $\delta$. Specifically, the rate improves as $\delta$ increases: for $0<\delta<2$, the decay follows ${\cal O}(k^{-\delta/2})$; at the critical case $\delta=2$, a logarithmic factor appears; and for $\delta>2$, the rate stabilizes at the optimal order ${\cal O}(1/k)$. These results demonstrate that, with appropriately chosen tolerances, the accelerated inexact PPA can approach the theoretically tight convergence behavior established in [6].
@@
Under the same regularity conditions as those used by Rockafellar [9], as discussed above, we now show that the accelerated version of the proximal point method still attains a fast linear convergence rate. Based on the results in Theorem 2.1, we can derive the following results in a manner similar to that of [10, Theorem 2], and thus, we provide only a sketch of the proof.
@@
Assume that Assumption 2.1 holds and $\{z^{k}\}$ is any sequence generated by the Halpern accelerated proximal point algorithm under criterion (B) of (2.2) with $c_{k}$ nondecreasing ($c_{k}\nearrow c_{\infty}\leq+\infty$). Assume that $\{z^{k}\}$ is bounded and $T^{-1}$ is Lipschitz continuous at 0 with modulus $a$; let
@@
The sequence $\{z^{k}\}$, being bounded, also satisfies criterion (A) of (2.2) for $\varepsilon_{k}=\delta_{k}\|\bar{z}^{k}-z^{k}\|$. Thus the conclusions of Theorem 2.1 hold. Using the same technique in [10, Theorem 2], we can obtain that, there exists a sufficient large number $\bar{k}$ such that
@@
4 Accelerated inexact augmented Lagrange method
Building upon the accelerated inexact proximal point method developed in the previous section, we now extend our analysis to its application in constrained convex optimization within the augmented Lagrangian framework. In analogy to Rockafellar’s seminal work [10, 9], the proposed accelerated inexact proximal point method naturally gives rise to an accelerated inexact augmented Lagrangian method. This connection enables us to establish convergence rate and iteration complexity results for the accelerated iALM in a manner parallel to the analysis of the classical inexact PPA. Throughout this section, we follow the notations in [9].
@@ 4 Accelerated inexact augmented Lagrange method
It follows from [9] that the augmented Lagrange method for solving (4.1) is the proximal point algorithm to the inclusion
@@
where $L$ is the augmented Lagrange function
@@
$\begin{array}[]{l}(C)\quad\quad\quad\quad\quad\quad\phi_{k}(x^{k})-\inf\phi_{k}\leq\varepsilon_{k}^{2}/2c_{k},\quad\displaystyle\sum_{k=0}^{\infty}\varepsilon_{k}<+\infty,\end{array}$ (4.10)
@@
Now we are in a position to state the accelerated augmented Lagrange method:
@@
As a direct consequence of [9, Proposition 6], we obtain the following result, which reveals the connection between criterion (C) for the inexact ALM and criterion (A) for the inexact PPM.
@@
Suppose $\sup(D)>$-$\infty$, and let the accelerated augmented Lagrange method (4.12) be executed. If the generated sequence $\{y^{k}\}\subset\mathbb{R}^{m}_{+}$ is bounded, then $y^{k}\rightarrow y^{\infty}$, where $y^{\infty}$ is some optimal solution to Problem (D), and $\{x^{k}\}$ is asymptotically minimizing for Problem (P) with
$f_{i}(x^{k+1})\leq\displaystyle\frac{1}{c_{k}}\left\{y^{k+1}_{i}-y^{k}_{i}+\displaystyle\frac{1}{k+1}[y^{k+1}_{i}-y^{0}_{i}]\right\}\rightarrow 0,\quad\mbox{ for }i=1,\ldots,m,$ (4.14)
$f_{0}(x^{k+1})-{\rm asym}\inf(P)\leq\displaystyle\frac{1}{2c_{k}}\left[\varepsilon_{k}^{2}+\|y^{k}\|^{2}-\|y^{k+1}\|^{2}-\displaystyle\frac{2k+3}{(k+1)^{2}}\left\langle y^{k+1}-y^{0},y^{k+1}-\displaystyle\frac{1}{2k+3}y^{0}\right\rangle\right].$ (4.15)
The boundedness of $\{y^{k}\}$ under (C) is actually equivalent to the existence of an optimal solution to Problem (D). It holds if Problem (P) satisfies the Slater condition, in this case one has $\max\,(D)=\inf\,(P)={\rm asym}\,\inf(P)$.
@@
Corollary 4.1 shows that criterion (C) implies criterion (A) for $T=-\partial g$. Thus the accelerated augmented Lagrange method is reduced to the Halpern accelerated proximal point algorithm under criterion (A) to $0\in\partial-g(y)$. Then from Theorem 2.1, we have that the sequence $\{y^{k}\}$ converges to a solution $y^{\infty}$ to $0\in-\partial g(y)$, namely an optimal solution to Problem (D). From the definition of $y^{k+1}$, we have
@@
Now we discuss the complexity of the accelerated augmented Lagrange method. Define the average point
@@
Suppose the solution set to Problem (D) is nonempty, and let the accelerated augmented Lagrange method (4.12) be executed. Then the generated sequences $\{x^{k}\}\subseteq C$ and $\{y^{k}\}\subset\mathbb{R}^{m}_{+}$ satisfy
@@
Suppose the solution set to Problem (D) is nonempty, and let the accelerated augmented Lagrange method (4.12) be executed in which $\varepsilon_{k}$ is defined by (3.20), and $\{x^{k}\}\subseteq C$ and $\{y^{k}\}\subset\mathbb{R}^{m}_{+}$ are generated sequences.
@@
$f_{i}(\tilde{x}^{k})\leq\displaystyle\left(\displaystyle\sum_{j=0}^{k-1}c_{j}\right)^{-1}\left\{3\Delta_{0}+\displaystyle\sum_{j=2}^{k}\frac{2\Delta_{0}}{j}\right\}\leq\left(kc_{0}\right)^{-1}\left\{3\Delta_{0}+\displaystyle\int_{1}^{k}\frac{2\Delta_{0}}{t}dt\right\}\leq\displaystyle\frac{1}{c_{0}k}[3\Delta_{0}+2\ln k],$

### v2

@@ 3 The rate of convergence
$\begin{array}[]{rl}0&\geq k(k+1)\|z^{k}-P_{c}(z^{k})\|^{2}+2(k+1)\langle z^{k}-P_{c}(z^{k}),z^{k}-z^{0}\rangle+(C_{1}+C_{2})+C_{3}+\displaystyle\sum_{j=1}^{k}\displaystyle\frac{2j^{2}}{j+1}\|\eta^{j-1}\|^{2}\\[6.0pt] &=k(k+1)\|z^{k}-P_{c}(z^{k})\|^{2}+2(k+1)\langle z^{k}-P_{c}(z^{k}),z^{k}-z^{0}\rangle+D_{1}+D_{2}+E_{1}+E_{2}+\displaystyle\sum_{j=1}^{k}\displaystyle\frac{2j^{2}}{j+1}\|\eta^{j-1}\|^{2}\\[6.0pt] &=k(k+1)\|z^{k}-P_{c}(z^{k})\|^{2}+2(k+1)\langle z^{k}-P_{c}(z^{k}),z^{k}-z^{0}\rangle+(D_{1}+E_{2})+(D_{2}+E_{1})+\displaystyle\sum_{j=1}^{k}\displaystyle\frac{2j^{2}}{j+1}\|\eta^{j-1}\|^{2}\\[6.0pt] &=k(k+1)\|z^{k}-P_{c}(z^{k})\|^{2}+2(k+1)\langle z^{k}-P_{c}(z^{k}),z^{k}-z^{0}\rangle-2k\left\langle z^{k}-z^{0},\eta^{k-1}\right\rangle\\[6.0pt] &\quad-\displaystyle\sum_{j=0}^{k-1}2\left\langle z^{0}-z^{j},\frac{j+1}{j+2}\eta^{j}\right\rangle-\displaystyle\sum_{j=0}^{k-1}2j\left\langle z^{0}-z^{j},\frac{j+1}{j+2}\eta^{j}-\eta^{j-1}\right\rangle-2k(k+1)\left\langle z^{k}-P_{c}(z^{k}),\eta^{k-1}\right\rangle\\[6.0pt] &\quad+2\displaystyle\sum_{j=0}^{k-1}\displaystyle\frac{(j+1)^{3}}{j+2}\left\langle P_{c}(z^{j})-z^{j},\eta^{j-1}-\eta^{j}\right\rangle-2\displaystyle\sum_{j=0}^{k-1}\displaystyle\frac{j+1}{j+2}\left\langle P_{c}(z^{j})-z^{j},\eta^{j-1}\right\rangle+\displaystyle\sum_{j=1}^{k}\displaystyle\frac{2j^{2}}{j+1}\|\eta^{j-1}\|^{2}.\end{array}$
@@
    $\sum_{j=0}^{k-1}\left\langle z^{0}-z^{j},\frac{j+1}{j+2}\eta^{j}\right\rangle\leq\displaystyle\sum_{j=0}^{k-1}\|z^{0}-z^{j}\|\cdot\|\frac{j+1}{j+2}\eta^{j}\|\leq\kappa_{0}\beta_{0}$;
@@
    $\sum_{j=0}^{k-1}\displaystyle\frac{j+1}{j+2}\left\langle P_{c}(z^{j})-z^{j},\eta^{j-1}\right\rangle\leq\displaystyle\sum_{j=0}^{k-1}\displaystyle\frac{j+1}{j+2}\|P_{c}(z^{j})-z^{j}\|\|\eta^{j-1}\|\leq\displaystyle\kappa_{0}\displaystyle\sum_{j=0}^{k-1}\displaystyle\varepsilon_{j-1}$.
@@
Next, we show that, with a specific choice of the tolerance, the Halpern accelerated proximal point method achieves a convergence rate that is very close to the tight rate established in [11].
@@
Corollary 3.1 provides explicit convergence rate estimates for the Halpern accelerated proximal point method under the tolerance rule (3.29). It shows that the residual $\|z^{k}-P_{c}(z^{k})\|$, which measures the violation of the monotone inclusion, converges to zero at a sublinear rate depending on the decay parameter $\delta$. Specifically, the rate improves as $\delta$ increases: for $0<\delta<2$, the decay follows ${\cal O}(k^{-\delta/2})$; at the critical case $\delta=2$, a logarithmic factor appears; and for $\delta>2$, the rate stabilizes at the optimal order ${\cal O}(1/k)$. These results demonstrate that, with appropriately chosen tolerances, the accelerated inexact PPM can approach the theoretically tight convergence behavior established in [11].
@@
Under the same regularity conditions as those used by Rockafellar [17], as discussed above, we now show that the accelerated version of the proximal point method still attains a fast linear convergence rate. Based on the results in Theorem 2.1, we can derive the following results in a manner similar to that of [18, Theorem 2], and thus, we provide only a sketch of the proof.
@@
Assume that Assumption 2.1 holds. Let $\{z^{k}\}$ and $\bar{z}^{k}$ be any sequences generated by the Halpern accelerated proximal point algorithm under criterion (B) of (2.4) with $c_{k}$ nondecreasing ($c_{k}\nearrow c_{\infty}\leq+\infty$). Assume that $\{z^{k}\}$ is bounded and $T^{-1}$ is Lipschitz continuous at 0 with modulus $a$; let
@@
The sequence $\{z^{k}\}$, being bounded, also satisfies criterion (A) of (2.4) for $\varepsilon_{k}=\delta_{k}\|\bar{z}^{k}-z^{k}\|$. Thus the conclusions of Theorem 2.1 hold. Using the same technique in [18, Theorem 2], we can obtain that, there exists a sufficient large number $\bar{k}$ such that
@@
4 Accelerated inexact augmented Lagrangian method
Building upon the accelerated inexact proximal point method developed in the previous section, we now extend our analysis to its application in constrained convex optimization within the augmented Lagrangian framework. In analogy to Rockafellar’s seminal work [18, 17], the proposed accelerated inexact proximal point method naturally gives rise to an accelerated inexact augmented Lagrangian method. This connection enables us to establish convergence rate and iteration complexity results for the accelerated iALM in a manner parallel to the analysis of the classical inexact PPM. Throughout this section, we follow the notations in [17].
@@ 4 Accelerated inexact augmented Lagrangian method
It follows from [17] that the augmented Lagrangian method for solving (4.37) is the proximal point algorithm to the inclusion
@@
where $L$ is the augmented Lagrangian function
@@
$\begin{array}[]{l}(C)\quad\quad\quad\quad\quad\quad\phi_{k}(x^{k+1})-\inf\phi_{k}\leq\varepsilon_{k}^{2}/2c_{k},\quad\displaystyle\sum_{k=0}^{\infty}\varepsilon_{k}<+\infty,\end{array}$ (4.46)
@@
Now we are in a position to state the accelerated augmented Lagrangian method:
@@
As a direct consequence of [17, Proposition 6], we obtain the following result, which reveals the connection between criterion (C) for the inexact ALM and criterion (A) for the inexact PPM.
@@
Suppose $\sup(D)>$-$\infty$, and let the accelerated augmented Lagrangian method (4.48) be executed. If the generated sequence $\{y^{k}\}\subset\mathbb{R}^{m}_{+}$ is bounded, then $y^{k}\rightarrow y^{\infty}$, where $y^{\infty}$ is some optimal solution to Problem (D), and $\{x^{k}\}$ is asymptotically minimizing for Problem (P) with
$f_{i}(x^{k+1})\leq\displaystyle\frac{1}{c_{k}}\left\{y^{k+1}_{i}-y^{k}_{i}+\displaystyle\frac{1}{k+1}[y^{k+1}_{i}-y^{0}_{i}]\right\}\rightarrow 0,\quad\mbox{ for }i=1,\ldots,m,$~~~~ (4.50)
$\begin{array}[]{l}f_{0}(x^{k+1})-{\rm asym}\inf(P)\\[5.69054pt] \leq\displaystyle\frac{1}{2c_{k}}\left[\varepsilon_{k}^{2}+\|y^{k}\|^{2}-\|y^{k+1}\|^{2}-\displaystyle\frac{2k+3}{(k+1)^{2}}\left\langle y^{k+1}-y^{0},y^{k+1}-\displaystyle\frac{1}{2k+3}y^{0}\right\rangle\right].\end{array}$ (4.51)
The boundedness of $\{y^{k}\}$ under (C) is actually equivalent to the existence of an optimal solution to Problem (D). It holds if Problem (P) satisfies the Slater condition; in this case one has $\max\,(D)=\inf\,(P)={\rm asym}\,\inf(P)$.
@@
Corollary 4.1 shows that criterion (C) implies criterion (A) for $T=-\partial g$. Thus the accelerated augmented Lagrangian method is reduced to the Halpern accelerated proximal point method under criterion (A) to $0\in\partial-g(y)$. Then from Theorem 2.1, we have that the sequence $\{y^{k}\}$ converges to a solution $y^{\infty}$ to $0\in-\partial g(y)$, namely an optimal solution to Problem (D). From the definition of $y^{k+1}$, we have
@@
Now we discuss the complexity of the accelerated augmented Lagrangian method. Define the average point
@@
Suppose the solution set to Problem (D) is nonempty, and let the accelerated augmented Lagrangian method (4.48) be executed. Then the generated sequences $\{x^{k}\}\subseteq C$ and $\{y^{k}\}\subset\mathbb{R}^{m}_{+}$ satisfy
@@
Suppose the solution set to Problem (D) is nonempty, and let the accelerated augmented Lagrangian method (4.48) be executed in which $\varepsilon_{k}$ is defined by (3.29), and $\{x^{k}\}\subseteq C$ and $\{y^{k}\}\subset\mathbb{R}^{m}_{+}$ are generated sequences.
@@
$f_{i}(\tilde{x}^{k})\leq\displaystyle\left(\displaystyle\sum_{j=0}^{k-1}c_{j}\right)^{-1}\left[3\Delta_{0}+\displaystyle\sum_{j=2}^{k}\frac{2\Delta_{0}}{j}\right]\leq\left(kc_{0}\right)^{-1}\left[3\Delta_{0}+\displaystyle\int_{1}^{k}\frac{2\Delta_{0}}{t}dt\right]\leq\displaystyle\frac{1}{c_{0}k}[3\Delta_{0}+2\ln k],$
@@
5 Numerical Experiments
In this section, we report numerical experiments designed to validate the theoretical results established in this paper. In particular, we investigate the performance of the inexact proximal point method with Halpern accelerated iteration (HiPPM) on two classes of optimization problems with distinct regularity properties. The first example is the nuclear norm regularized least squares problem [14], which does not necessarily satisfy the regularity assumptions required for the linear convergence analysis. The second example is the sparse logistic regression problem, which satisfies metric subregularity and hence admits a fast linear convergence rate for the inexact proximal point method even in the absence of Halpern acceleration. All experiments are implemented in Matlab 2018b.
5.1 Nuclear norm regularized Least square problem
The following nuclear norm regularized least squares problem is commonly used as a convex surrogate for the rank minimization problem:
$\min_{z\in\mathbb{R}^{m\times n}}\;\frac{1}{2}\|Az-B\|^{2}+\lambda\|z\|_{*},$ (5.68)
where $\lambda>0$ and $\|z\|_{*}$ denotes the nuclear norm of $z$, defined as the sum of its singular values.
In this experiment, we adopt the stopping criterion (A) with $\varepsilon_{k}$ specified by (3.29) to validate the main result stated in Theorem 3.1. Recall that the stopping criterion is given by
$\|\bar{z}^{k}-P_{c}(z^{k})\|\leq\varepsilon_{k}=\frac{1}{(k+2)^{1+\delta}}.$ (5.69)
Here, $\bar{z}^{k}$ is an approximate solution obtained by solving the $k$ th subproblem of the HiPPM,
$\min_{z\in\mathbb{R}^{m\times n}}\;\Phi_{k}(z):=\frac{1}{2}\|Az-B\|^{2}+\lambda\|z\|_{*}+\frac{1}{2c}\|z-z^{k}\|^{2}.$ (5.70)
We solve the subproblem (5.70) using a dual-based accelerated gradient descent (AGD) method. Specifically, AGD is applied to the corresponding dual problem
$\max_{x}\;\Psi_{k}(x):=-\frac{1}{2}\|x\|^{2}-\langle B-Az^{k},x\rangle+c^{-1}\mathcal{M}_{c\varphi}(z^{k}-cA^{T}x)-\frac{c}{2}\|A^{T}x\|^{2},$ (5.71)
where $\mathcal{M}_{c\varphi}(\cdot)$ denotes the Moreau envelope of $\varphi(z):=\lambda\|z\|_{*}$.
Following the discussion in [22], the stopping criterion involving the unknown proximal point $P_{c}(Z^{k})$ can be replaced by the following computable condition:
$\|\bar{z}^{k}-P_{c}(z^{k})\|\leq\sqrt{2c\bigl(\Phi_{k}(\bar{z}^{k})-\Psi_{k}(x^{k+1})\bigr)}\leq\varepsilon_{k},$ (5.72)
where $x^{k+1}$ is an approximate solution to the dual problem (5.71), and
$\bar{z}^{k}=cP_{c}\bigl(c^{-1}z^{k}-A^{T}x^{k+1}\bigr).$
In the experiment, we set $m=300$ and $n=200$, and generate a ground truth matrix with rank $r=50$. The matrices $A\in\mathbb{R}^{50\times 300}$ and $B\in\mathbb{R}^{50\times 200}$ are generated randomly. The regularization parameter is chosen as $\lambda=1$, and the proximal parameter is fixed at $c=10$. Since the Halpern iteration is sensitive to the choice of the anchor point $Z^{0}$, the selection of the initial point plays a critical role in the convergence behavior. To mitigate this effect, we adopt a heuristic restarting strategy similar to that used in [9]. All reported results are averaged over 20 independent runs.
We first compare the performance of the iPPM, the HiPPM without restarting, and the HiPPM with restarting every 10 and 20 iterations, respectively. The performance is evaluated in terms of the KKT residual, measured on a base-10 logarithmic scale, defined as
$r(z)=\log_{10}\!\left(\|z-P_{c}(z-A^{T}(Az-B))\|\right).$
The numerical results are summarized in Figure 1. To highlight the effect of the Halpern accelerate iteration, we present a magnified view of the first 20 iterations. As shown in the figure, the HiPPM exhibits faster convergence than the iPPM during the early iterations. However, as the number of iterations increases, the convergence speed of the HiPPM without restarting may deteriorate due to the influence of the fixed anchor point. In contrast, the restarting strategy effectively alleviates this issue and enables the HiPPM to consistently achieve a faster convergence rate than the iPPM.
Figure 1: Comparison of the KKT residuals for iPPM and HiPPM with and without restarting.
Since Theorem 3.1 characterizes the convergence behavior of the quantity $\|Z^{k}-P_{c}(Z^{k})\|$ under the prescribed stopping criterion (5.72), we first note that
$\|z^{k}-P_{c}(z^{k})\|\leq\|z^{k}-\bar{z}^{k}\|+\|\bar{z}^{k}-P_{c}(z^{k})\|\leq\|z^{k}-\bar{z}^{k}\|+\varepsilon_{k}.$
Therefore, under the adopted stopping rule, it suffices to investigate the convergence rate of $\|z^{k}-\bar{z}^{k}\|$. We compare its behavior for different values of $\delta\in\{1,2,3\}$ against the baseline ${\rm log}_{10}(\alpha/(k+1))$, where $\alpha$ is defined as the average of $\|z^{0}-z^{*}\|$ over all runs, and $z^{*}$ denotes an approximate solution obtained when the stopping condition r(z)<6r(z)<{-6} is satisfied in each run. The corresponding results are reported in Figure 2. The numerical results show that, in the absence of restarting, the decay rate is initially slower than the baseline, and the case $\delta=3$ is close to the baseline. As the iteration count increases, the convergence behavior for all three values of $\delta$ becomes comparable to that of the baseline. Moreover, the restarting strategy is observed to significantly accelerate convergence.
Figure 2: Decay of $\|z^{k}-\overline{z}^{k}\|$ for different values of $\delta\in\{1,2,3\}$, compared with the baseline $\alpha/(k+1)$.
5.2 Fast linear convergence rate under regularity conditions
We consider the following sparse logistic regression problem:
$\min_{z\in\mathbb{R}^{n}}\;h(Az)+\lambda\|z\|_{1},$ (5.73)
where, for a given vector $b\in\mathbb{R}^{m}$, the function $h:\mathbb{R}^{m}\to\mathbb{R}$ is defined as
$h(y)=\sum_{i=1}^{m}\log\bigl(1+e^{-b_{i}y_{i}}\bigr),\qquad\forall\,y\in\mathbb{R}^{m}.$
Let $T(z):=\partial\bigl(h(Az)+\lambda\|z\|_{1}\bigr)$ denote the associated maximal monotone operator. According to [10, Theorem 3.3], the operator $T$ satisfies an error bound condition at a point 0 with modulus $a\geq 0$. Specifically, there exists a constant $\bar{\varepsilon}>0$ such that, for any $z\in\mathbb{R}^{n}$ satisfying ${\rm dist}(0,T(x))\leq\bar{\varepsilon}$, one has
${\rm dist}\bigl(z,T^{-1}(0)\bigr)\leq a\,{\rm dist}\bigl(0,T(x)\bigr).$
Moreover, it follows from [10, Theorem 3.2] that, under the above regularity condition, the convergence results established in Theorem 4.1 remain valid.
In this experiment, we evaluate the fast linear convergence behavior predicted by Theorem 4.1 using two benchmark binary classification datasets from the UCI repository, namely colon-cancer and duke. The colon-cancer dataset consists of 62 samples with 2000 features, while the duke dataset contains 44 samples with 7129 features. We set the regularization parameter to $\lambda=10^{-4}$ and incorporate the Halpern iteration into the implementation based on the publicly available code [^4] https://github.com/linmeixia/exclusive-lasso-solver. The subproblems are solved using a semismooth Newton method, and the proximal parameter $c_{k}$ is chosen to increase progressively along the iterations.
We report the evolution of the quantity $\alpha_{k}:={\|\bar{z}^{k}-\bar{z}\|}/{\|z^{k}-\bar{z}\|}$ for both datasets under different restarting frequencies in Figure 3. The numerical results show that $\alpha_{k}<1$ for all iterations, which implies the parameter $\vartheta_{k}$ in (3.35) is less than $1$. Furthermore, as the parameter $c_{k}$ increases, the values of $\alpha_{k}$ decrease accordingly and thus confirms the fast linear convergence behavior, which is consistent with the theoretical predictions of Theorem 3.2.
Figure 3: Solving problem with regularity condition, $\alpha_{k}={\|\bar{z}^{k}-\bar{z}\|}/{\|z^{k}-\bar{z}\|}$
6 Conclusion
In this paper, we studied a Halpern-type acceleration of the inexact proximal point method for maximal monotone inclusion problems in Hilbert spaces. We established its global convergence and developed a unified framework for convergence rate analysis, showing that the squared fixed-point residual converges at an $\mathcal{O}(1/k^{2})$ rate under mild inexactness conditions. Moreover, we proved that the proposed Halpern-accelerated scheme achieves fast linear convergence rate under some regularity assumption. We further extended the analysis to constrained convex optimization via the augmented Lagrangian framework, deriving convergence rate and complexity results for the resulting accelerated inexact augmented Lagrangian method. From a practical perspective, the numerical performance can be further improved by restarting strategies. However, the restart mechanism adopted in this work is heuristic and the design of theoretically justified and adaptive restarting rules remains an important direction for future research.